# UGotMe: An Embodied System for Affective Human-Robot Interaction

### v1

@@ Abstract
Equipping humanoid robots with the capability to understand emotional states of human interactants and express emotions appropriately according to situations is essential for affective human-robot interaction. However, enabling current vision-aware multimodal emotion recognition models for affective human-robot interaction in the real-world raises embodiment challenges: addressing the environmental noise issue and meeting real-time requirements. First, in multiparty conversation scenarios, the noises inherited in the visual observation of the robot, which may come from either 1) distracting objects in the scene or 2) inactive speakers appearing in the field of view of the robot, hinder the models from extracting emotional cues from vision inputs. Secondly, real-time response, a desired feature for an interactive system, is also challenging to achieve. To tackle both challenges, we introduce an affective human-robot interaction system called UGotMe designed specifically for multiparty conversations. Two denoising strategies are proposed and incorporated into the system to solve the first issue. Specifically, to filter out distracting objects in the scene, we propose extracting face images of the speakers from the raw images and introduce a customized active face extraction strategy to rule out inactive speakers. As for the second issue, we employ efficient data transmission from the robot to the local server to improve real-time response capability. We deploy UGotMe on a human robot named Ameca to validate its real-time inference capabilities in practical scenarios. Videos demonstrating real-world deployment are available at https://pi3-141592653.github.io/UGotMe/
@@ I Introduction
In healthcare and service environments, humanoid robots are increasingly engaging with people, particularly through conversational interactions. To enhance the naturalness and fluidity of human-robot conversations, it is essential to endow these robots with the capability to perceive and understand human emotional states and to express emotions in a manner that resembles human behavior. However, enabling current vision-aware multimodal emotion recognition models for affective human-robot interaction raises challenges: addressing the environmental noise issue and meeting real-time requirements.
@@
In particular, emotion recognition module of the system, the proposed Vision-Language to Emotion (VL2E) model, is designed to be compatible with the aforementioned denoising strategies, i.e., the vision encoder of VL2E consumes face sequences and extracts features at frame level. It models intra-modal interactions within visual features using self-attention transformer. In addition, to ensure the robot can understand human emotions in a manner similar to human interaction, we incorporate context modeling by considering the most recent $k$ utterances in the dialogue when calculating textual representations. We employ multimodal transformer to fuse features from visual and textual modalities. On the MELD dataset, VL2E outperforms all baselines in weighted average F1. An illustration of VL2E is shown in Fig. 3. We aim to enable humanoid robots to convey emotions via facial expressions, as they are recognized as a highly effective means of communicating affective information compared to verbal cues and tone [2]. Therefore, upon emotion recognition, we directly map them to robotic facial expressions in line with parallel empathy (generating the same emotion as the peer) [3]. In the present study, we only have access to robotic facial expressions within a predefined set, corresponding to seven basic emotions (including neutral, surprise, fear, sadness, joy, disgust, and anger). These expressions can be applied to the humanoid robot named Ameca, which was developed by Engineered Arts. We deploy UGotMe on Ameca, demonstrating its real-time inference capabilities in practical scenarios.
@@ III-C Emotion Recognition in Conversation
Inspired by prompt learning [27], we append a prompt after the conversation context for $t$-th turn as follows:
@@
respectively. In the present study, we use the SimCSE [28] model as our text encoder.
@@ IV-C Evaluation
For the embodied system, UGotMe, the evaluation metrics include emotion response accuracy and user experience (rated on a scale from 0 to 10, with higher scores indicating better performance). Both metrics are assessed by human raters with four raters included in this study. Emotion response accuracy is first calculated within each dialogue and then averaged across dialogues for each rater. The final scores, averaged across all raters, will be reported
@@ V Results
(a) Ameca responds with an angry face when the active speaker said “Dylon cheated me again, I do no wanna talk to him anymore” in anger.
@@ V-C Real-world Deployment
Models for emotion recognition are hosted on a local server with one NVIDIA RTX 4090 GPU. Videos can be viewed at https://pi3-141592653.github.io/UGotMe/.

### v2

@@ Abstract
Equipping humanoid robots with the capability to understand emotional states of human interactants and express emotions appropriately according to situations is essential for affective human-robot interaction. However, enabling current vision-aware multimodal emotion recognition models for affective human-robot interaction in the real-world raises embodiment challenges: addressing the environmental noise issue and meeting real-time requirements. First, in multiparty conversation scenarios, the noises inherited in the visual observation of the robot, which may come from either 1) distracting objects in the scene or 2) inactive speakers appearing in the field of view of the robot, hinder the models from extracting emotional cues from vision inputs. Secondly, real-time response, a desired feature for an interactive system, is also challenging to achieve. To tackle both challenges, we introduce an affective human-robot interaction system called UGotMe designed specifically for multiparty conversations. Two denoising strategies are proposed and incorporated into the system to solve the first issue. Specifically, to filter out distracting objects in the scene, we propose extracting face images of the speakers from the raw images and introduce a customized active face extraction strategy to rule out inactive speakers. As for the second issue, we employ efficient data transmission from the robot to the local server to improve real-time response capability. We deploy UGotMe on a human robot named Ameca to validate its real-time inference capabilities in practical scenarios. Videos demonstrating real-world deployment are available at https://lipzh5.github.io/HumanoidVLE/
@@ I Introduction
Humanoid robots and AI [1] are increasingly engaging with people, particularly through conversational interactions, for applications like healthcare and services. To enhance the naturalness and fluidity of human-robot conversations, it is essential to endow these robots with the capability to perceive and understand human emotional states and to express emotions in a manner that resembles human behavior. However, enabling current vision-aware multimodal emotion recognition models for affective human-robot interaction raises challenges: addressing the environmental noise issue and meeting real-time requirements.
@@
In particular, emotion recognition module of the system, the proposed Vision-Language to Emotion (VL2E) model, is designed to be compatible with the aforementioned denoising strategies, i.e., the vision encoder of VL2E consumes face sequences and extracts features at the frame level. It models intra-modal interactions within visual features using self-attention transformer. In addition, to ensure the robot can understand human emotions in a manner similar to human interaction, we incorporate context modeling by considering the most recent $k$ utterances in the dialogue when calculating textual representations. We employ multimodal transformer to fuse features from visual and textual modalities. On the MELD dataset, VL2E outperforms all baselines in weighted average F1. An illustration of VL2E is shown in Fig. 3. We aim to enable humanoid robots to convey emotions via facial expressions, as they are recognized as a highly effective means of communicating affective information compared to verbal cues and tone [3]. Therefore, upon emotion recognition, we directly map them to robotic facial expressions in line with parallel empathy (generating the same emotion as the peer) [4]. In the present study, we only have access to robotic facial expressions within a predefined set, corresponding to seven basic emotions (including neutral, surprise, fear, sadness, joy, disgust, and anger). These expressions can be applied to the humanoid robot named Ameca, which was developed by Engineered Arts. We deploy UGotMe on Ameca, demonstrating its real-time inference capabilities in practical scenarios.
@@ III-C Emotion Recognition in Conversation
To help the model distinguish between the context and the target turn, we adopt an approach inspired by prompt learning [28], appending a prompt $P_{t}$ to the conversation context for the $t$-th turn. The prompt $P_{t}$ is defined as follows:
@@
respectively.
@@ IV-C Evaluation
For the embodied system, UGotMe, the evaluation metrics include emotion response accuracy and user experience (rated on a scale from 0 to 10, with higher scores indicating better performance). Both metrics are assessed by human raters with four raters included in this study. Emotion response accuracy is first calculated within each dialogue and then averaged across dialogues for each rater. The final scores, averaged across all raters, will be reported.
@@ V Results
(a) Ameca responds with an angry face when the active speaker said “Dylon cheated me again, I don’t wanna talk to him anymore” in anger.
@@ V-C Real-world Deployment
Models for emotion recognition are hosted on a local server with one NVIDIA RTX 4090 GPU. Videos can be viewed at https://lipzh5.github.io/HumanoidVLE/.
@@ VI Conclusion
VII ACKNOWLEDGMENT
This work was partly supported by Australian Research Council Discovery Grant DP240102050, LIEF Grant LE240100131 and Linkage Grant LP230201022.